# Modulating CNN Features with Pre-Trained ViT Representations for Open-Vocabulary Object Detection

### v1

@@ Abstract
Owing to large-scale image-text contrastive training, pre-trained vision language model (VLM) like CLIP shows superior open-vocabulary recognition ability. Most existing open-vocabulary object detectors attempt to utilize the pre-trained VLM to attain generative representation. F-ViT uses the pre-trained visual encoder as the backbone network and freezes it during training. However, the frozen backbone doesn’t benefit from the labeled data to strengthen the representation. Therefore, we propose a novel two-branch backbone network design, named as ViT-Feature-Modulated Multi-Scale Convolutional network (VMCNet). VMCNet consists of a trainable convolutional branch, a frozen pre-trained ViT branch and a feature modulation module. The trainable CNN branch could be optimized with labeled data while the frozen pre-trained ViT branch could keep the representation ability derived from large-scale pre-training. Then, the proposed feature modulation module could modulate the multi-scale CNN features with the representations from ViT branch. With the proposed mixed structure, detector is more likely to discover novel categories. Evaluated on two popular benchmarks, our method boosts the detection performance on novel category and outperforms the baseline. On OV-COCO, the proposed method achieves 44.3 AP with ViT-B/16 and 48.5 AP with ViT-L/14. On OV-LVIS, VMCNet with ViT-B/16 and ViT-L/14 reaches 27.8 and 38.4 mAPr.
@@ 1 Introduction
With the advances in cross-modality learning, open-vocabulary object detection [32] (OVOD) is proposed. This task requires models to detect the novel targets beyond the training-category set via the language instruction. Compared to traditional closed-set detector, OVOD model improves the generalization ability from instance-level to category-level, which makes a further step in the real-world applications.
@@
For most existing OVOD methods, pre-trained VLM is essential to implement open-vocabulary recognition. CLIP [18], a typical VLM pre-trained on an enormous number of image-text pairs, is able to encode image into the feature aligned with the text embedding space. To better satisfy the dense prediction task such as object detection, some works [37, 36, 27, 7] further improve CLIP, which transfer the visual encoder‘s representation ability from image-level to region-level. Based on the dense visual feature encoder, F-ViT [27] firstly builds open-vocabulary object detectors upon a frozen CLIP ViT backbone and achieves remark performance. In F-ViT, its backbone network is fine-tuned exclusively in a self-distillation manner at first. During the base training, F-ViT freezes its backbone to save the generalization ability. As shown in Figure 1(a), F-ViT uses the interpolated intermediate features from the frozen CLIP ViT for detection.
@@
It is common knowledge that backbone network plays a key role in detection performance. In most detector designs, the backbone networks are optimized with the labeled bounding boxes during training, which ensures the detection performance. However, to save the generalization ability, the backbone network of F-ViT gives up exploiting the knowledge of labeled data during base training. Although its self-distillation strategy strengthens the local representation of CLIP ViT which may implicitly promotes detection, this pretext task treats the vision transformer more like a region classifier than a backbone network. Backbone network, as the feature extractor, needs to suit the downstream modules in a detector via the bounding box training. Thus, merely using self-distillation may not be adequate for feature extraction of detection.
@@
A usual scheme in object detection is to initialize the backbone with the weights from pre-training and fine-tune it with training data. However, directly fine-tuning the pre-trained model with limited data will harm its open-vocabulary recognition ability, which degrades the model into a closed-set detector. As reported in [27], the performance on novel categories will degenerate when using trainable CLIP ViT. Besides, another straightforward scheme is applying an extra trainable network to extract features parallelly for detection as shown in Figure 1(b). Whereas the additional network exploits the training data, such design isolates the CLIP ViT, which does not make use of the representation ability of VLM for detection. Hence, it is nature to ask a question: could a backbone structure utilize the information from both the base training data and the pre-trained VLM simultaneously?
@@
Inspired by ViT-CoMer [30], we are motivated to design a two-branch backbone network, which could take advantages of the base training and pre-trained CLIP ViT. The proposed backbone network, named as ViT-Feature-Modulated Multi-Scale Convolutional network (VMCNet), consists of a vision transformer branch, a convolutional branch and a feature modulation module. Figure 1(c) shows the sketch of our design. Firstly, to retain the generalization ability, we also freeze the CLIP ViT but use it as a part of network. These ViT features are treated as intermediate products instead of the final output of backbone. Then, to exploit the base training, we use a trainable convolutional neural network (CNN) as the convolutional branch. This CNN branch has a simple structure and is able to produce multi-scale features. Finally, the proposed feature modulation module fuses the outputs from two branches to generate the final representation. As a result, VMCNet improves the detection performance on novel categories of OVOD detector.
@@
We propose a novel backbone network for open-vocabulary object detection, named as ViT-Feature-Modulated Multi-Scale Convolutional network (VMCNet). Via utilizing two-branch architecture, it leverages the strengths of different networks and exploits the information from both the base training data and the pre-trained VLM.
We design a feature modulation module to merge the features from different branches. By this module, multi-scale features from trainable convolutional network are modulated with the adapted feature from frozen ViT. The process of feature fusion improves the quality of representation and boosts the detection performance on novel categories.
We evaluate our proposed VMCNet on two popular benchmarks. The experimental results demonstrate that our method significantly improves the detect performance on novel categories. On OV-COCO, our method with ViT-B/16 achieves 44.3 AP. When utilizing ViT-L/14, VMCNet could achieve 48.5AP. Moreover, VMCNet with ViT-B/16 and ViT-L/14 attains 27.8 and 38.4 mAPr on OV-LVIS, which is competitive with SOTA methods.
@@ 2 Related Work
Backbone Structure for Object Detection. In deep vision tasks, deep neural network was firstly applied in image classification. The advent of convolutional neural network (CNN) such as ResNet [9] strongly promoted the development of vision tasks. R-CNN[3] started to use the neural network pre-trained on classification dataset for feature extraction. From then on, employing backbone network in detector became a common practice. Besides CNN, transformer [22] architecture was introduced into visual field. In ViT [2], an image is divided into patches then encoded into a sequence. These attention-based designs [14, 24, 34] also exhibited superior performances in vision tasks . Recently , [39] proposes a backbone network based on state space models (SSMs), which serves as a new option for detector design. To further exploit the strength of different architectures, some works [5, 15, 30] design the hybrid approaches which further boost the dense visual prediction tasks.
@@
Inspired by ViT-CoMer [30], our design chooses the mixed-architecture which combines ViT and CNN. Distinct from these closed-set detection designs, not all modules in our backbone network are trainable. To satisfy the open-vocabulary setting, our backbone network attempts to exploit the information not only from the training data, but also the pre-trained VLM. Thus, a frozen CLIP ViT is employed as a branch of backbone network, and our feature fusion module performs unidirectional information injection instead of bidirectional interaction.
Backbone design for Open-Vocabulary Object Detection. One of the major issues in OVOD is how to alleviate over-fitting. So far, most OVOD designs still rely on bounding box training to enable the localization ability. However, this attribute inevitably leads to over-fitting on base categories. Another fact is that most OVOD detectors inherit the model structures from the closed-set designs. Backbone network, as the indispensable part of detector, faces the conflict between capturing localization ability and alleviate over-fitting on base-data.
Many works such as [4, 26, 23, 17, 35] choose to optimize the backbone networks with base data for detection, especially these based on the two-stage detectors like [8]. A common strategy is to replace the class-specific module into the class-agnostic and train the entire detector. Usually, the architecture of these backbones is CNN-based. With the additional improvement such as knowledge distillation [4, 23], pseudo labels [35] and etc, the detectors could perform well for OVOD task.
There are also some works [37, 37, 29] using the frozen backbone networks. Although ViT visual encoder outperforms the ResNet in image-level recognition, the CNN-based network can preserve region attribute better for dense prediction. With modification on the final layer like [37], CLIP ResNet could produce the dense feature for detection. Following two-stage detector structure, F-VLM [12] freezes its backbone and fine-tunes only the detector head during base training. CORA [29], the query-based method, also extracts image features with the frozen CLIP ResNet. To explore the potential of CLIP ViT on dense prediction task such as detection, CLIPSelf utilizes a self-distillation strategy, which enhances the recognition ability of CLIP ViT on region-level. Then, F-ViT uses this frozen CLIP ViT as the backbone network to realize the OVOD.
Unlike these single-branch backbone designs, our method utilizes the two-branch structure which consists of both the trainable branch and the frozen. Two types of features could be efficiently fused into a stronger representation.
@@ 3.1 Preliminary
Our detector structure is based on F-ViT. Before detailing our backbone design, we will give a brief description of F-ViT to get a clear view of the entire OVOD method. Unless specified, we choose ViT-B/16 as the default setting for illustration.
The detection task could be divided into two sub-tasks, i.e., localization and classification. F-ViT follows the two-stage detector design [8]. The RoI features for classification and localization regression are derived from the intermediate features from the frozen ViT. We denote the output features of th transformer block in CLIP ViT as , prediction of the detector could be represented as:
@@
Since the region-level representation of ViT is enhanced with self-distillation, F-ViT also applies the dense feature from last transformer block to improve the classification prediction. Specifically, they use RoI Align to attain the region feature with RoI bounding box . After normalization, the VLM score is attained via calculating the cosine similarity between and the encoded text feature . The final classification score is the weighted geometric average of and , where and are hyper-parameters.
In our design, we focus on improving the quality of by using the proposed backbone network. Except the backbone design, we follow the detector design of F-ViT to implement open-vocabulary object detection.
@@ 3.2 Overall Architecture
Our goal is to design a backbone which could efficiently utilize the information from both the base training data and pre-trained vision transformer. The overall architecture of VMCNet is illustrated in Figure 2, which mainly includes three parts: (a) Multi-Scale Convolutional Branch. (b) Vision Transformer Branch. (c) ViT-Feature-Modulated Multi-Scale Convolutional(VMC) Module. These three components will be introduced specifically in Section 3.3, 3.4 and 3.5, respectively.
The input image passes through both branches parallelly to obtain the multi-scale convolutional features and vision transformer features. Then, two types of features are fed into Vit-Modulated Convolutional (VMC) module, which modulates the multi-scale CNN features with the adapted ViT representation. The output of VMC module is then processed via batch normalization and transpose convolution to attain the final feature . could serve as the backbone features for downstream processing. Note that contains the features with resolutions of 1/4, 1/8, 1/16 and 1/32, which is same as the setting in F-ViT.
@@ 3.3 Multi-Scale Convolutional Branch
Convolutional neural network is well-known for its properties of local continuity and multi-scale capabilities, besides, it usually have less computation burden than the attention-based network. Thus, we choose to utilize CNN as the optimizable branch to learn detection representation from base training. Inspired by ViT-CoMer [30], we utilize the early stage of its convolutional branch to extract multi-scale convolutional features. As shown in Figure 2, this light-weight version of CNN branch consists of a stem module [9], a stack of convolutional layers, layer embedding addition and a multi-receptive field feature pyramid (MRFP) module [30]. By using such structure, the additional computation cost is acceptable and we can customize the channel number of output.
Firstly, stem module processes an input image with the shape of into the feature with a resolution reduction of 1/4 of the original image. Then, this intermediate feature goes through three stride-2 convolutional layers to obtain the features with resolutions of 1/8, 1/16, and 1/32, respectively.
Next, feature at each scale is projected by its corresponding convolutional layer. Except feature at the largest scale, projected features are added with layer embeddings. The output features are denoted as , , , and , respectively. For saving the computation cost, outputs excluding are fed into MRFP and VMC.
The structure of MRFP is specified in [30], this module could refine multi-scale features efficiently and expand receptive field. , and are flattened and concatenated into . We study the choice of MRFP number in Section 4.3.1. The output of convolutional branch is attained as follow, where :
@@ 3.4 Vision Transformer Branch
Pre-trained on an enormous number of image-text pairs, visual encoder in CLIP[18] could process an image into the encoded feature aligned with text feature space. CLIPSelf [27] further improved the visual transformer encoder with self-distillation for dense prediction task. Thus, we use the frozen visual encoder from [27] as ViT branch to attain the modulating feature.
In ViT branch, input image firstly passes through a large kernel convolutional layer (patch embedding). After added with the positional embedding, we attain the input feature . is fed into a series of vision transformer blocks to attain its intermediate features.
To make full use of the representation ability of the frozen ViT, we collect multiple intermediate features from this branch. Under ViT-B/16 setting, we choose , and for feature modulation. Study about the choice of ViT layers is presented in Section 4.3.2. Since not all transformer blocks are used during training phase, we could skip the computation of remaining blocks during training. In inference phase, we apply the whole vision transformer to generate feature map from the last block for dense prediction.

### v2

@@ Abstract
Owing to large-scale image-text contrastive training, pre-trained vision language model (VLM) like CLIP shows superior open-vocabulary recognition ability. Most existing open-vocabulary object detectors attempt to utilize the pre-trained VLMs to attain generalized representation. F-ViT uses the pre-trained visual encoder as the backbone network and freezes it during training. However, its frozen backbone doesn’t benefit from the labeled data to strengthen the representation for detection. Therefore, we propose a novel two-branch backbone network, named as ViT-Feature-Modulated Multi-Scale Convolutional Network (VMCNet), which consists of a trainable convolutional branch, a frozen pre-trained ViT branch and a VMC module. The trainable CNN branch could be optimized with labeled data while the frozen pre-trained ViT branch could keep the representation ability derived from large-scale pre-training. Then, the proposed VMC module could modulate the multi-scale CNN features with the representations from ViT branch. With this proposed mixed structure, the detector is more likely to discover objects of novel categories. Evaluated on two popular benchmarks, our method boosts the detection performance on novel category and outperforms state-of-the-art methods. On OV-COCO, the proposed method achieves 44.3 AP with ViT-B/16 and 48.5 AP with ViT-L/14. On OV-LVIS, VMCNet with ViT-B/16 and ViT-L/14 reaches 27.8 and 38.4 mAPr.
@@ 1 Introduction
Open-vocabulary object detection [34] (OVOD) requires models to detect the novel targets beyond the training-category set via the language instruction. Compared to traditional closed-set detector, OVOD model improves the generalization ability from instance-level to category-level, which makes a further step in the real-world applications such as autonomous driving, intelligent androids, medical imaging analysis, etc.
@@
For most existing OVOD methods, pre-trained VLM is essential to implement open-vocabulary recognition. CLIP [19], a typical VLM pre-trained on an enormous number of image-text pairs, is able to encode an image into the feature aligned with the text embedding space. To better satisfy the dense prediction task such as object detection, some works [39, 38, 29, 7] further improve CLIP, which transfer the visual encoder’s representation ability from image-level to region-level. Based on the dense visual feature encoder, F-ViT [29] firstly builds an open-vocabulary object detector upon a frozen CLIP ViT backbone and achieves remarkable performance. Specifically, the ViT from CLIP is fine-tuned in a self-distillation manner at first. Then, F-ViT freezes this ViT during base training and leverages it as the backbone network to support detection. Sketch of the feature flow in F-ViT is shown in fig. 1(a).
@@
It is common knowledge that backbone network plays a key role in detection performance. In most detector designs, the backbone networks are optimized with the labeled bounding boxes during training, which ensures the detection performance. However, to save the generalization ability, the backbone network of F-ViT fails to exploit the knowledge of labeled data during base training. Although its self-distillation strategy strengthens the local representation of CLIP ViT which may implicitly promote detection, this strategy treats the vision transformer more like a region classifier than a backbone network. As the feature extractor, backbone network needs to suit the downstream modules in a detector via the bounding box training. Thus, merely using self-distillation may not be adequate for feature extraction of detection.
@@
A usual scheme in object detection is to initialize the backbone with the weights from pre-training and fine-tune it with training data. However, directly fine-tuning the pre-trained model with limited data will harm its open-vocabulary recognition ability, which degrades the model into a closed-set detector. As reported in [29], the performance on novel categories will degenerate when using trainable ViT backbone. Another straightforward scheme is applying an extra trainable network to extract features parallelly for detection as shown in fig. 1(b). Although the additional network exploits the training data, such design isolates the CLIP ViT, which does not leverage the representation ability of VLM for detection. Hence, it is nature to ask a question: could a backbone structure utilize the information from both the base training data and the pre-trained VLM simultaneously?
@@
Inspired by ViT-CoMer [32], we propose a two-branch backbone network, which could make use of the base training and pre-trained CLIP ViT. The proposed backbone network, named as ViT-Feature-Modulated Multi-Scale Convolutional Network (VMCNet), consists of a vision transformer branch, a multi-scale convolutional branch and a VMC module. fig. 1(c) shows the sketch of our design. Firstly, to exploit the base training, we use a trainable convolutional neural network (CNN) as the convolutional branch. This CNN branch has a simple structure and is able to produce multi-scale features. Optimized with labeled bounding box, this branch could learn from base training data and provide multi-scale detection features. Besides, we need the representations from pre-trained VLM to strengthen the generalization ability of features. To reach this goal, we also freeze the CLIP ViT and use it as the second branch in our network. These ViT features are treated as intermediate products instead of the final output of backbone. Finally, given the features from two branches, the proposed VMC module modulates the multi-scale CNN features with the adaptive representations from ViT. By using the proposed structure, two branches are complementary to each other, and the knowledge from both the base training data and pre-trained ViT could be exploited to generate the detection features. As a result, VMCNet improves the detection performance of OVOD detector on novel categories.
@@
We propose a novel backbone network for open-vocabulary object detector, named as ViT-Feature-Modulated Multi-Scale Convolutional Network (VMCNet). Via utilizing two-branch architecture, it leverages the strengths of different networks and exploits the information from both the base training data and the pre-trained VLM.
We propose the VMC module to fuse the features from different branches. By this module, multi-scale features from trainable convolutional network are modulated with the adapted features from frozen ViT. This feature fusion processing improves the quality of representation and boosts the detection performance on novel categories.
We evaluate our proposed VMCNet on two popular benchmarks. The experimental results demonstrate that our method significantly improves the detection performance on novel categories. On OV-COCO, our method with ViT-B/16 achieves 44.3 AP. When utilizing ViT-L/14, VMCNet could achieve 48.5 AP. Moreover, VMCNet with ViT-B/16 and ViT-L/14 attains 27.8 and 38.4 mAPr on OV-LVIS, which surpasses state-of-the-art methods.
@@ 2 Related Work
Backbone Structure for Object Detection. In deep vision tasks, deep neural network was firstly applied in image classification. Typical CNNs such as AlexNet [12], VGGNet [21] and ResNet [9] strongly promoted the development of vision tasks. R-CNN[3] started to use the neural network pre-trained on classification dataset for feature extraction. From then on, employing backbone network in detector became a common practice. Besides CNN, transformer [24] architecture was introduced into vision field. In ViT [2], an image is divided into patches and then encoded into a sequence. These attention-based networks [15, 26, 36] also exhibited superior performances in vision tasks. Recently, Vim [41] proposes a backbone network based on state space models (SSMs), which serves as a new option for detector design. To exploit the strength of different architectures, some works [5, 16, 32] devise the hybrid approaches which further boost the dense visual prediction tasks.
@@
Inspired by ViT-CoMer [32], our method chooses the mixed-architecture which combines ViT and CNN. Distinct from these closed-set detection designs, not all modules in our backbone network are trainable. To satisfy the open-vocabulary setting, our backbone network attempts to exploit the information not only from the training data, but also the pre-trained VLM. Thus, a frozen CLIP ViT is employed as a branch of the proposed network, and our feature fusion module performs unidirectional information injection instead of bidirectional interaction.
Backbone design for Open-Vocabulary Object Detection. One of the major issues in OVOD is how to alleviate over-fitting. So far, most OVOD methods still rely on bounding box training to enable the localization ability. However, this attribute inevitably leads to over-fitting on base categories. Another fact is that most OVOD detectors inherit the model structures from the closed-set designs. Backbone network, as the indispensable part of detector, faces the conflict between capturing localization ability and alleviating over-fitting on base-data.
Many works such as [4, 28, 25, 18, 37] choose to optimize the backbone networks with base data for detection, especially these based on the two-stage detectors like [8]. A common strategy is to replace the class-specific module into the class-agnostic and train the entire detector. Usually, the architecture of their backbones is CNN-based. With the additional improvement such as knowledge distillation [4, 25], pseudo labels [37], etc., the detectors could perform well for OVOD task.
There are also some works [13, 31, 29] using the frozen backbone networks. Although ViT visual encoder outperforms the ResNet in image-level recognition, the CNN-based network can preserve region attribute better for dense prediction. With modification on the final layer like [39], CLIP ResNet could produce the dense feature for detection. Following two-stage detector structure, F-VLM [13] freezes its backbone and fine-tunes only the detector head during base training. CORA [31], the query-based method, also extracts image features with the frozen CLIP ResNet. To explore the potential of CLIP ViT on dense prediction task, CLIPSelf [29] utilizes a self-distillation strategy, which enhances the recognition ability of CLIP ViT on region-level. Then, F-ViT [29] uses this frozen CLIP ViT as the backbone network to realize the OVOD.
Unlike these single-branch backbone designs, our method utilizes the two-branch structure which consists of both the trainable CNN and the frozen ViT. Two types of features could be efficiently fused into a stronger representation by the proposed VMC module.
@@ 3.1 Preliminary
Our detector structure is based on F-ViT [29]. Before detailing our backbone design, we will give a brief description of F-ViT to get a clear view of the entire OVOD method. Unless specified, we choose ViT-B/16 as the default setting for illustration.
The detection task could be divided into two sub-tasks, i.e., localization and classification. F-ViT follows the two-stage detector design [8]. The RoI features for classification and localization regression are derived from the interpolated intermediate features from the frozen ViT. We denote the output features of -th transformer block in CLIP ViT as , prediction of the detector could be represented as:
@@
Since the region-level representation of ViT is enhanced with self-distillation, F-ViT also applies the dense feature from last transformer block to improve the classification prediction. Specifically, they use RoI Align to attain the region feature with RoI bounding box . After normalization, the VLM score is attained via calculating the cosine similarity between and the encoded text features . The final classification score is the weighted geometric average of and , where and are hyper-parameters.
In our method, we focus on improving the quality of by using the proposed backbone network. Except the backbone part, we follow the detector design of F-ViT to implement open-vocabulary object detection.
@@ 3.2 Overall Architecture
Our goal is to design a backbone which could efficiently utilize the information from both the base training data and pre-trained vision transformer. The overall architecture of VMCNet is illustrated in fig. 2, which mainly includes three parts: (a) Multi-Scale Convolutional Branch. (b) Vision Transformer Branch. (c) ViT-Feature-Modulated Multi-Scale Convolutional (VMC) Module. These three components will be introduced specifically in section 3.3, section 3.4 and section 3.5, respectively.
The input image passes through both branches parallelly to obtain the multi-scale convolutional features and vision transformer features. Then, two types of features are fed into VMC module, which modulates the multi-scale CNN features with the adapted ViT representations. The output of VMC module is then processed via batch normalization and transposed convolution to attain . serves as the backbone features and will be sent to downstream processing. Note that contains the features with resolutions of 1/4, 1/8, 1/16 and 1/32, which is same as the setting in F-ViT.
@@ 3.3 Multi-Scale Convolutional Branch
Convolutional neural network is well-known for its properties of local continuity and multi-scale capabilities. Besides, it usually have less computation burden than the attention-based network. Thus, we choose to utilize CNN as the optimizable branch to learn detection representation from base training. Inspired by ViT-CoMer [32], we utilize the early stage of its convolutional branch to extract multi-scale convolutional features. As shown in fig. 2, this light-weight version of CNN branch consists of a stem module [9], a stack of convolutional layers, level embedding addition and a multi-receptive field feature pyramid (MRFP) module [32]. By using such structure, the additional computation cost is acceptable and we can customize the channel number of output.
Firstly, stem module processes an input image with the shape of into the feature with a resolution reduction of 1/4 of the original image. Then, this intermediate feature goes through three 2-stride convolutional layers to obtain the features with resolutions of 1/8, 1/16, and 1/32, respectively.
Next, feature at each scale is projected by its corresponding convolutional layer. Except feature at the largest scale, projected features are added with level embeddings. The output features are denoted as , , , and , respectively. For saving the computation cost, outputs excluding are fed into MRFP and VMC module.
MRFP [32] could refine multi-scale features efficiently and expand receptive field. Further details about MRFP may be found in [32]. , and are flattened and concatenated into . We study the choice of MRFP number in section 4.3.1. The output of convolutional branch is attained as follows, where :
@@ 3.4 Vision Transformer Branch
Pre-trained on an enormous number of image-text pairs, visual encoder in CLIP [19] could process an image into the encoded feature aligned with text feature space. CLIPSelf [29] further improves the visual transformer encoder with self-distillation for dense prediction task. Thus, we use the frozen visual encoder from [29] as ViT branch to attain the modulating feature.
In ViT branch, input image firstly passes through a large kernel convolutional layer (patch embedding). After added with the positional embedding, we attain the input feature . is fed into a series of vision transformer blocks to attain the intermediate features.
To make full use of the representation ability of the frozen ViT, we collect multiple intermediate features from this branch. Under ViT-B/16 setting, we choose , and for feature modulation. Study about the choice of ViT layers is presented in section 4.3.2. Since not all transformer blocks are used to attain the intermediate features, we could skip the computation of remaining blocks during training. In inference phase, we apply the whole vision transformer to generate feature map from the last block for dense prediction.